# Predictive Performance Comparison of Decision Policies Under Confounding

### v1

@@ 6.2 Doubly robust estimator
Therefore, we propose a doubly robust approach for estimating $\hat{v}$ that corrects the bias in the plug-in estimator. Because doubly robust estimators have an error term which is a product of nuisance function errors, they attain fast convergence rates, even when estimating nuisance functions at slow non-parametric rates (Kennedy et al., 2021). For example, to attain $n^{-1/2}$ rates for the estimator, it is sufficient to estimate nuisance functions at $n^{-1/4}$, a rate achieved by many non-parametric machine learning techniques. The form of the doubly robust estimator depends on the pointwise bounding functions implied by a causal assumption.
@@
Algorithm 1 (Appendix C) can be leveraged for doubly-robust estimation of policy regret bounds by substituting e.q. 2 for estimation of $\mathcal{H}(\hat{v}_{1,k}(t,0);\hat{\tau})$. We show in Appendix C that this estimator has second order error in nuisance estimation error.
@@ 7 Numerical experiments
where $\sigma(x)=\frac{1}{1+e^{-x}}$ and $\sigma^{\prime}(x)$ is the softmax function. We use Algorithm 1 with the plug-in estimator to learn estimates estimates of the $\delta$-regret interval, where we fit $\hat{\eta}$ via a non-parametric gradient boosting classifier. We compare against the baseline regret interval by applying bounds provided in Appendix A over the same $\hat{\mathcal{V}}(p,\hat{\tau})$ used to estimate the $\delta$-regret. See Appendix D for additional setup details.
Figure 5 provides a comparison of the $\delta$-regret interval and baseline regret interval across five policy performance measures. In line with our theoretical results, we obtain tighter bounds from the $\delta$-regret across all policy performance measures apart from the positive predictive value (PPV). We show the oracle regret evaluated via $Y(1)$ in black. In Figure 5, we plot coverage of the estimated $\delta$-regret interval evaluated against the true population value, denoted by dashed lines. The bottom panel shows $95\%$ bootstrap confidence intervals around the true upper and lower $\delta$-regret interval. These results show that estimates concentrate around the true regret interval as the number of samples increases. Appendix D contains additional experimental results stress-testing coverage under causal assumption violations.
@@ Appendix B Assumption mapping extensions and proofs
In this appendix, we discuss additional causal assumptions which imply uncertainty sets over partially-identified $v$-statistics. Rosenbaum’s $\Gamma$-sensitivity model (Rosenbaum, 2005), the proximal identification framework (Ghassami et al., 2023), and Manski-style no assumptions bounds (Manski & Pepper, 1998) imply bounding functions $\tau(\cdot)$, which can be used to construct $\mathcal{V}(p;\tau)$ by invoking Lemma 5.2.
@@ B.5 Manski style no-assumptions bounds
Manski-style no assumptions bounds on partially identified policy comparison terms (Manski, 1989; Manski & Pepper, 1998) follow by setting $\underline{\tau}(x)=0$, $\overline{\tau}(x)=1$. Invoking Lemma 5.2 yields $\mathcal{H}(v_{y}(t,0);\tau)=[0,p(D=0,T=1)]$, which is the same interval recovered by the unconstrained uncertainty set $\mathcal{V}(p;\tau)$.
@@ B.6 Assumption mapping proofs
Appendix C Estimation Algorithm and Theoretical Results
Algorithm 1 Plug-in regret bound estimator
  for $k=1$ to $K$ do
     Learn $\hat{\eta}_{k}=(\hat{e}_{1,k},\hat{\tau}_{k})$ using $\mathcal{O}_{-k}$
     Construct $\hat{\mathcal{V}}_{k}(p;\hat{\tau})$
  end for
C.1 Proof of Theorem 6.1
@@ Appendix C Estimation Algorithm and Theoretical Results
$P(\hat{f}-f)$ $=P\left(\hat{e}(X)\cdot\hat{\tau}(X)-e(X)\cdot\tau(X)\right)$ (10)
$=P\left(\hat{e}(X)\cdot\hat{\tau}(X)+\hat{e}(X)\cdot\tau(X)-\hat{e}(X)\cdot\tau(X)-e(X)\cdot\tau(X)\right)$ (11)
$=P\left(\hat{e}(X)(\hat{\tau}(X)-\tau(X))+\tau(X)\cdot(\hat{e}(X)-e(X))\right)$ (12)
$=P\left(\hat{e}(X)(\hat{\tau}(X)-\tau(X)))+P(\tau(X)\cdot(\hat{e}(X)-e(X))\right)$ (13)
$\leq||\hat{e}(x)-e(x)||+||\hat{\tau}(x)-\tau(x)||$ (14)
@@
C.2 Rate of doubly robust estimator
@@ D.0.4 Real-world data experiments
Setup Details. Each record of the synthetic dataset provided by (Obermeyer et al., 2019) contains patient demographics, health information measured in the prior year, and outcome variables measured in the following year. Only $\approx 1\%$ of patients were enrolled in the care management program under the status quo policy. This low selection rate yields vacuous regret intervals without strong assumptions on confounding. Therefore, for this illustration, we construct a status quo policy with an $\approx 18\%$ selection rate ($N=2452$) by including all 452 records marked as enrolled in the program and randomly sampling 2000 unenrolled records. We fit the cost prediction model and nuisance functions using patient demographic variables, comorbidities, and prescription information measured in the preceding year. We use a linear regression to predict patient cost, and threshold predictions at the 55th percentile. This cutoff matches the threshold for physician enrollment recommendations of the deployed risk assessment (Obermeyer et al., 2019). We use $40\%$ of the subsampled data to fit the cost prediction model and the remaining $60\%$ as an observational sample for the policy comparison task. Because we do not have access to physician identifiers which can be used as an instrument, we leverage the MSM assumption for identification. We use a gradient boosting classifier to fit nuisance functions. We use Algorithm 1 to estimate $[\underline{\hat{R}}_{\delta}(\pi,\pi_{0};m,\mathcal{\hat{V}}),\hat{\overline{R}}_{\delta}(\pi,\pi_{0};m,\mathcal{\hat{V}})]$ under the MSM with with $\Lambda=1.2$. We report results over $N=20$ trials with $K=2$ folds.
@@ E.1 Sources of uncertainty impacting policy comparisons
Confounding-related uncertainty: We can understand the impact of confounding related uncertainty on off-policy evaluation via the target regression decomposition
$\mathbb{E}[Y(1)\mid X=x]=\underbrace{\mathbb{E}[Y(1)\mid D^{\pi_{0}}=1,X=x]}_{\text{Identified from observational data}}\cdot p(D^{\pi_{0}}=1\mid X=x)+\underbrace{\mathbb{E}[Y(1)\mid D^{\pi_{0}}=0,X=x]}_{\in[0,1],\text{Unidentified from observational data}}\cdot p(D^{\pi_{0}}=0\mid X=x).$

### v2

@@ 6.2 Doubly Robust Estimator
Therefore, we propose a doubly robust approach for estimating $\hat{v}$ that corrects the bias in the plug-in estimator. Because doubly robust estimators have an error term which is a product of nuisance function errors, they attain fast convergence rates, even when estimating nuisance functions at slow non-parametric rates (Kennedy et al., 2023). For example, to attain $n^{-1/2}$ rates for the estimator, it is sufficient to estimate nuisance functions at $n^{-1/4}$, a rate achieved by many non-parametric machine learning techniques. The form of the doubly robust estimator depends on the pointwise bounding functions implied by a causal assumption.
@@
Algorithm 1 can be leveraged for doubly-robust estimation of policy regret bounds by substituting e.q. 2 for estimation of $\mathcal{H}(\hat{v}_{1,k}(t,0);\hat{\tau})$. We show in Appendix C that this estimator has second order error in nuisance estimation error.
@@ 7 Numerical Experiments
where $\sigma(x)=\frac{1}{1+e^{-x}}$ and $\sigma^{\prime}(x)$ is the softmax function. We use Algorithm 1 with the plug-in estimator to learn estimates estimates of the $\delta$-regret interval. We compare against the baseline regret interval by applying bounds provided in Appendix A over the same $\hat{\mathcal{V}}(p,\hat{\tau})$ used to estimate the $\delta$-regret. See Appendix D for additional setup details.
Figure 5 provides a comparison of the $\delta$-regret interval and baseline regret interval across five policy performance measures. In line with our theoretical results, we obtain tighter bounds from the $\delta$-regret across all policy performance measures apart from the positive predictive value (PPV). We show the oracle regret evaluated via $Y(1)$ in black. In Figure 5, we plot coverage of the estimated $\delta$-regret interval evaluated against the true population value, denoted by dashed lines. The bottom panel shows $95\%$ bootstrap confidence intervals around the true upper and lower $\delta$-regret and baseline regret intervals. These results show that estimates concentrate around the true regret interval as the number of samples increases. We observe that the baseline interval yields better coverage of the oracle regret in small sample settings because it is more conservative and thus more tolerant to bias around estimates of the asymptotic interval end points. Appendix D contains additional experimental results stress-testing coverage under causal assumption violations.
@@ 9 Conclusion
Acknowledgements
We thank our anonymous reviewers, the attendees of the NeurIPS 2023 Workshop on Regulatable Machine Learning, and the attendees of Carnegie Mellon University’s Fairness, Explainability, Accountability, and Transparency (FEAT) reading group for their helpful feedback. This work was supported by an award from the UL Research Institutes through the Center for Advancing Safety of Machine Intelligence (CASMI) at Northwestern University and the National Science Foundation Graduate Research Fellowship Program (Award No. DGE1745016).
@@ Appendix B Assumption Mapping Extensions and Proofs
In this appendix, we discuss additional causal assumptions which imply uncertainty sets over partially-identified $v$-statistics. Rosenbaum’s $\Gamma$-sensitivity model (Rosenbaum, 2005), the proximal identification framework (Ghassami et al., 2023), and Manski-style no assumptions bounds (Manski & Pepper, 2000) imply bounding functions $\tau(\cdot)$, which can be used to construct $\mathcal{V}(p;\tau)$ by invoking Lemma 5.2.
@@ B.5 Manski Style No-Assumptions Bounds
Manski-style no assumptions bounds on partially identified policy comparison terms (Manski, 1989; Manski & Pepper, 2000) follow by setting $\underline{\tau}(x)=0$, $\overline{\tau}(x)=1$. Invoking Lemma 5.2 yields $\mathcal{H}(v_{y}(t,0);\tau)=[0,p(D=0,T=1)]$, which is the same interval recovered by the unconstrained uncertainty set $\mathcal{V}(p;\tau)$.
@@ B.6 Assumption Mapping Proofs
Appendix C Theoretical Estimation Results
C.0.1 Proof of Theorem 6.1
@@ Appendix C Theoretical Estimation Results
$P(\hat{f}-f)$ $=P\left(\hat{e}(X)\cdot\hat{\tau}(X)-e(X)\cdot\tau(X)\right)$
$=P\left(\hat{e}(X)\cdot\hat{\tau}(X)+\hat{e}(X)\cdot\tau(X)-\hat{e}(X)\cdot\tau(X)-e(X)\cdot\tau(X)\right)$
$=P\left(\hat{e}(X)(\hat{\tau}(X)-\tau(X))+\tau(X)\cdot(\hat{e}(X)-e(X))\right)$
$=P\left(\hat{e}(X)(\hat{\tau}(X)-\tau(X)))+P(\tau(X)\cdot(\hat{e}(X)-e(X))\right)$
$\leq||\hat{e}(x)-e(x)||+||\hat{\tau}(x)-\tau(x)||$
@@
C.1 Rate of doubly robust estimator
@@ D.0.4 Real-world data experiments
Setup Details. Each record of the synthetic dataset provided by (Obermeyer et al., 2019) contains patient demographics, health information measured in the prior year, and outcome variables measured in the following year. Only $\approx 1\%$ of patients were enrolled in the care management program under the status quo policy. This low selection rate yields vacuous regret intervals without strong assumptions on confounding. Therefore, for this illustration, we construct a status quo policy with an $\approx 18\%$ selection rate ($N=2452$) by including all 452 records marked as enrolled in the program and randomly sampling 2000 unenrolled records. We fit the cost prediction model and nuisance functions using patient demographic variables, comorbidities, and prescription information measured in the preceding year. We use a linear regression to predict patient cost, and threshold predictions at the 55th percentile. This cutoff matches the threshold for physician enrollment recommendations of the deployed risk assessment (Obermeyer et al., 2019). We use $40\%$ of the subsampled data to fit the cost prediction model and the remaining $60\%$ as an observational sample for the policy comparison task. Because we do not have access to physician identifiers which can be used as an instrument, we leverage the MSM assumption for identification. We use a logistic regression classifier to fit nuisance functions. We use Algorithm 1 to estimate $[\underline{\hat{R}}_{\delta}(\pi,\pi_{0};m,\mathcal{\hat{V}}),\hat{\overline{R}}_{\delta}(\pi,\pi_{0};m,\mathcal{\hat{V}})]$ under the MSM with with $\Lambda=1.2$. We report results over $N=20$ trials with $K=2$ folds.
@@ E.1 Sources of uncertainty impacting policy comparisons
Confounding-related uncertainty: Let $\mu(x)=\mathbb{E}[Y(1)\mid X=x]$ be the target outcome regression function. We can understand the impact of confounding related uncertainty on off-policy evaluation via the target regression decomposition
$\mu(x)=\underbrace{\mathbb{E}[Y(1)\mid D^{\pi_{0}}=1,X=x]}_{\text{Identified from observational data}}\cdot p(D^{\pi_{0}}=1\mid X=x)+\underbrace{\mathbb{E}[Y(1)\mid D^{\pi_{0}}=0,X=x]}_{\in[0,1],\text{Unidentified from observational data}}\cdot p(D^{\pi_{0}}=0\mid X=x).$